### v2

@@ II.3.3 Thin outer layer
If we look at the temporal development of the magnetic energy and its spatial distribution, we recognize two different types of dynamo action. If the wall diffusivity is close to that of the fluid (i.e. ), a dynamo only occurs at very large (). This dynamo is time-independent (no oscillatory components, see orange and green curve in Figure 4a) and the associated field energy is essentially concentrated in a narrow region near the end caps (see iso-surfaces in Figure 4c). If the outer layer has an larger than (see blue curve in Figure 4b), we obtain an oscillating solution that already emerges at much smaller , which is exemplified by the blue curve for in Figure 4a. In that case the magnetic field energy is widely distributed throughout the entire volume (however still with a characteristic pattern, see Figure 4d.) The impact of the outer layer is illustrated by the growth rates shown in Figure 4b. When the magnetic diffusivity of the outer layer is increased to emulate the effect of a stainless steel container, an eigenmode is excited, the growth rates of which are represented by the inverted parabola in the individual curves in Figure 4b. This special shape results in positive growth rates for and for further increasing diffusivities the behaviour increasingly approaches that of pseudo-vaccum boundary conditions without outer layer (see black curve). It is therefore the eigenmode that belongs to the inverse parabolic growth rate, which corresponds to a dynamo that occurs at relatively low achievable in the experiment, and causes a magnetic field that permeates the entire volume. Interestingly, this eigenmode belongs to an oscillating dynamo while the much less efficient dynamo is stationary, although the opposite behavior is generally assumed [30].
II.3.4 Thick outer layer
@@
The beneficial effect of the outer layer with sufficiently large diffusivity vanishes, when its thickness is increased, as can be recognized in Figure 5a which shows the evolution of the magnetic energy for while the magnetic diffusivity remains at . Here we end up with a critical magnetic Reynolds number slightly above and the solution is again characterized by a non-oscillatory behaviour with the magnetic energy constraint to two small section close to the endcaps, which clearly indicates that we are in the low efficiency branch (see Figure 5c). Obviously, the dominant solution in this case is again given by the less efficient branch.
@@
Nevertheless, the highly efficient branch can be achieved again by assuming a more realistic configuration, where the container wall is emulated by a thin co-rotating layer with , while the laboratory is modeled by a thick outer layer with large diffusivity (actually, this should be an insulator in order to emulate air). In this case, we obtain an oscillating solution that occupies the entire cylinder volume and has a similar structure as in the case of the thin layer or pseudo-vacuum boundary conditions without an outer layer (see Figures 5b and 5d). However, the critical magnetic Reynolds number of is then about two and a half times as large as in these cases.
Thus we see from the kinematic linear models that there are two different branches of dynamo solutions, the realization of which depends on the configuration of the outer boundary regions, i.e. essentially on the boundary conditions. We have not investigated here at which exact parameters the transition from one branch to the other takes place, since the experimental realization must in any case start from a basically infinitely large outer volume of an insulator, namely the air in the laboratory. Interestingly, the resulting critical magnetic Reynolds number of is very close to the value derived by Goepfert and Tilgner [20] for the precession dynamo with the help of simulations in a cube.
In the following, we now want to solve the induction equation coupled to the Navier-Stokes equation with a comparable setup in a self-consistent MHD model, so that the time evolution of magnetic field and velocity field is solved simultaneously, including the feedback via the Lorentz force.
III Full set of magnetohydrodynamic equations
III.1 Code scheme
@@
An extension of the kinematic problem arises when we additionally consider the temporal evolution of the flow field. In this case, the Navier-Stokes equation needs to be solved as well, which in the case of precession-driven flow takes the form
@@
In addition to the terms commonly encountered in regular flow problems describing nonlinear interactions and dissipation, we have three contributions: the Coriolis force due to rotation, the Lorentz force due to the interaction with the magnetic field, and the Poincaré force accounting for acceleration due to perpetual change of the orientation of the cylinder’s axis of rotation. A detailed derivation of the Navier-Stokes equation for a precessing system is given e.g. in Tilgner [17].
@@
We conduct direct numerical simulations (DNS) of the coupled system of equations (1) and (2) using the code SpecDyn[14, 31], which is based on a pseudo-spectral Fourier approach. The drawback of this approach is the requirement of periodic boundary conditions in all three directions. Since this rules out an application of a cylindrical geometry, the scheme is complemented by a direct forcing immersed boundary method (IBM) which allows the consideration of nearly arbitrary shaped boundaries that prescribe the confinement for the fluid flow when the boundary conditions on these boundaries are of no-slip type. The code is parallelized using a slab decomposition and a standard MPI library. Since the most important part of the calculations results from the computation of the Fast Fourier Transformation (FFT), special efforts are undertaken to apply an extremely efficient scheme The fulfilment of the divergence-free condition for the velocity field is ensured using a standard Pressure-Corrector method where the pressure serves as a Lagrangian multiplier so that is ensured (i.e. first, solve Equation (2) without pressure, then project the intermediate solution on the divergence-free space by calculating the necessary pressure field). The same scheme is also applied to the magnetic field to ensure that the solution of the induction equation is also divergence-free as required by Maxwell’s equations. The code has been tested with various configurations that proved accurate solutions with excellent convergence properties and nearly perfect strong scaling properties. A detailed description is given in Wilbert, Giesecke, and Grauer [14] and a comparable approach was used for numerical models of the VKS dynamo[32, 33].
III.2 Model set-up
@@
We consider a set-up given by a cylindrical fluid region with radius and height embedded in a Cartesian domain with a edge length so that the size of the box provides sufficient space to let the magnetic field decline outside the cylindrical vessel. The cylindrical vessel is implemented by imposing internal conditions on the fluid that emulate the (virtual) container walls by enforcing no-slip conditions at the surface of the container applying the direct forcing immersed boundary method and we do not model any special region representing the wall of the vessel because of such a layer would not be properly resolved. Instead we use a magnetic diffusivity for the fluid region, whereas the ’non-liquid’ region outside the vessel has a magnetic diffusivity , which is supposed to emulate an electrically isolating exteriour. The simulations are carried out with a Reynolds number of which is the maximum possible at the resolution of that still sufficiently resolves the boundary layers at the fluid-wall interface of the vessel [31].
III.3 Results
@@
Similar to the procedure in the kinematic case, the Poincaré number is varied, whereby different magnetic Reynolds numbers are considered, beginning with high values. If a positive mean growth rate is found for a particular set of parameters, further runs are executed consecutively lowering the value of until no more dynamo action can be observed. Rather similar to the kinematic models we find a limited range for the preferred occurrence of a dynamo in the interval . However, the critical magnetic Reynolds number is significantly higher with a minimum value of for . The slightly larger value for the optimum compared to the kinematic models can be explained by the (slightly) smaller (see also [15]). In the following we focus on the particular case and and compare the properties of this DNS solution with the kinematic approximation obtained at and . This paradigmatic examples well characterize the solution and allow a faster growth with shorter time period until a statistically steady state, while showing similar qualitative behavior as the dynamos found very close to the threshold. Here we focus on this particular and compare the properties of this DNS solution with and in the kinematic approximation.
III.3.1 Temporal evolution of the kinetic energy
@@
The temporal evolutions of kinetic and magnetic energy are displayed in Fig. 6, where the kinetic and the magnetic energy are defined as
@@
The simulations start with the pure hydrodynamic case and the effects by magnetic induction are switched on at . After a short initial phase the magnetic energy begins to grow for about cylinder rotations and finally goes into saturation. The induced magnetic energy is much smaller than and both energies are dominated by the contributions from the components. We find that, even in the saturated state, the presence of the magnetic field does not substantially alter the structure and the amplitude of the fluid flow via the Lorentz force. A striking property is that unlike the kinetic energy, which remains relatively steady, the magnetic energy is characterized by quasi-periodic peaks where its value suddenly rises to times above the average.
Although there are no substantial changes in the flow field due to the back-reaction of the magnetic field, the bursts can be clearly assigned to corresponding peaks in the interaction parameter , which compares the amplitude of the Lorentz force to the inertial term in the Navier-Stokes equation:
which in case of a dynamo follows Miralles, Plihon, and Pinton [34].
III.3.2 Dynamo bursts and field structure
The peaks in the magneitc energy are always associated with a very strong concentration of magnetic energy in the form of local patches, as can be seen in the three-dimensional representation (see Fig 8). These patches are short-lived despite the strong concentration of magnetic energy, and even if they appear regularly, no corresponding period can be assigned. The field structure obtained from the simulations is visualized in Figure 8 which shows nested isosurfaces of the magnetic energy density for a couple of typical snapshots. We find a small scale field in terms of seemingly randomly distributed in space and time. The magnetic field patches are slightly elongated in the azimuthal direction (which goes along with the decrease of the scale in the perpendicular direction) and in general we find a spatial scale of the order of the energy carrying turbulence. The magnetic field patches exhibit an intermittent character and the thickness of the structures at least in one dimension is determined by the resistive scale. Due to the special structure of the Poincare force and the associated mean flow in the form of a standing forced inertial wave, we are not dealing with an isotropic, homogeneous turbulence, so that Kazantsev’s theory (see e.g. [brandenburgXY]) is not directly applicable.
Let us label the second (major) peak as case A, the third peak as case B, and the non-peak solution as case C (see Fig. 6). The magnetic field is small-scale in all cases, whereas the velocity field is large-scale. Although our self-exciting dynamo is small-scale, ongoing hydromagnetic processes tend to excite larger scales (i.e., a large-scale magnetic field), as seen in Figure 8a. and 8b. However, these large-scale magnetic field structures are quite isolated; they do not extend throughout the entire volume and quickly dissipate due to strong magnetic diffusion, typically within 1-2 rotation periods.
In contrast, the velocity field is large-scale, characterized by two vortex structures near the cylinder wall, and is concentrated mainly near the cylinder boundary, while it is almost zero in the central volume due to the dominance of the component. The vorticity, , naturally follows the velocity characteristics, being concentrated at the cylinder’s border and negligible toward the central volume, as shown in Fig. LABEL:B-U-O. The velocity field remains largely unchanged in cases A and B, retaining characteristics similar to those in case C. The vorticity, however, penetrates slightly deeper into the central volume compared to case C (see Fig. LABEL:B-U-O), which represents the standard off-peak solution, where the magnetic field is weak and small-scale without any significant large-scale structures.
@@ III.3.2 Dynamo bursts and field structure
While the distribution of the magnetic energy in individual snapshots appears to be largely chaotic (with the exception of the bursts), this is no longer the case when we look at the time average in the turntable system. With sufficiently long averaging of more than 200 rotation periods, we get the distribution shown in Figure 9.
III.3.3 Backreaction and kinetic power spectrum
A reduction in the speed amplitude is also not absolutely necessary, because the dynamo that sets in is a small-scale dynamo. For this case, theoretical considerations have shown that (in a mean-field model) the feedback effect via the Lorentz force corresponds to an increased diffusivity or hyperdiffusivity [35]. However, an reaction of the fluid flow can be seen in a subtle way when looking at the temporal power-spectra of kinetic energy. Figure 10 compares the temporal Power-spectrum of the kinetic energy for the hydrodynamic case (left panel) with the spectrum for the dynamo case (right panel). We see that, in the hydrodynamic case, the contributions to high frequencies by the axial component (spectral index ) drop significantly faster than the contributions of the horizontal components and (spectral index ).
In the dynamo case, the spectral index of the horizontal components approaches which is close to the axial component, which practically does not change compared to the hydrodynamic case. This change is not particularly strong, but it is still clearly noticeable.
IV Conclusions
@@
The kinematic models reveal the existence of two different branches for dynamo action. One branch is characterized by a highly localized and stationary dynamo that only occurs at rather large magnetic Reynolds number. The second type of dynamo is of oscillatory type with the magnetic field filling a larger proportion of the fluid domain. This type can occur at much lower magnetic Reynolds number, which, however requires suitable boundary conditions in terms of either a thin outer layer with large magnetic diffusivity or a comparable emulation with pseudo-vacuum BC. However, the kinematic model, which we assume best reflects the experimental setup, has a critical magnetic Reynolds number of , which is about twice as high as achievable in the planned precession dynamo experiment. Interestingly, this agrees quite well with the value derived by Goepfert and Tilgner [20] by means of simulations in a precessing cube.
@@
We then applied direct numerical simulations and have calculated a comparative solution in order to investigate the time evolution of the velocity field and how it is influenced by the magnetic field through the Lorentz force and to check for consistentency with the kinematic models. The configuration of this DNS essentially corresponds to that of the kinematic model, in which a stationary outer layer with large magnetic diffusivity emulates the outer space in the laboratory. Unfortunately, it is not possible to achieve the large Reynolds numbers, as they are relevant in the experiment, due to the limitations of the numerics and the hardware, respectively. This is mainly due to the extremely thin boundary layers, which require a sufficiently high resolution. As a consequence, we find dynamo solutions in the DNS only for relatively large magnetic Reynolds numbers so that we effectively deal with a magnetic Prandtl number of the order of one with the magnetic field being small-scale and fluctuating on short time scales. At this stage it cannot be concluded with certainty that this behavior will change if we move to larger Reynolds numbers, as there are currently insufficient parameter studies available for such scaling. For the parameters considered, a small-scale burst dynamo is the typical solution. These bursts are characterized by peaks in magnetic energy, which, as previously mentioned, represent the sudden excitation of large-scale magnetic field structures. Since these structures are isolated and not sustained throughout the entire volume, they are rapidly destroyed by strong magnetic diffusion. This is followed by a period without the excitation of large-scale structures (no peaks in magnetic energy) until the hydromagnetic dynamo once again excites some large-scale magnetic field structures. Our results align with findings [36, 37, 38], which conclude that increasing the magnetic Prandtl number leads to a transition from large-scale dynamos to small-scale dynamos. Thus, at high values of , small-scale dynamos predominate, which is consistent with our case.
@@
To summarize, the characteristics of the magnetic fields from the kinematic simulations are clearly different from those of the DNS. However, if the magnetic field from the DNS is averaged over a sufficiently long period of time so that the occasional occurrence of the strongly localized patches is also averaged out, the structure is comparable to the structure of the weakly efficient branch in the kinematic models, which only appears at large magnetic Reynolds numbers. This explains, at least in part, the high magnetic Reynolds number required to obtain a dynamo in the DNS.
Averaging the magnetic fields over a sufficient period of time, the small scale magnetic field structures give a similar pattern as for the weak branch of the kinematic solutions, while the occasional bursts appear to have a similar character to the strong branch, but fail to establish a corresponding dynamo solution.

### v3

@@ III.3 Wall layer
A detailed view on the impact of the diffusivity of the outer layer on the growthrates is shown in Figure 6. When the magnetic diffusivity of the outer layer is increased to emulate the effect of a stainless steel container, an eigenmode is excited, the growth rates of which are represented by the inverted parabola in the individual colored curves in Figure 6. For this special shape results in positive growth rates at far below the point where the "regular" curve becomes positive. For further increasing diffusivities the behaviour approaches that of pseudo-vacuum boundary conditions without outer layer (illustrated by the black curve). It is therefore the eigenmode that belongs to the inverse parabolic growth rate, which corresponds to a dynamo that occurs at relatively low and causes a magnetic field that permeates the entire volume.
@@
The beneficial effect of the wall layer with (sufficiently) large vanishes, when its thickness is increased, as shown in Figure 7a which presents the evolution of the magnetic energy for while the magnetic diffusivity remains at . Here we end up with a critical magnetic Reynolds number slightly above and the solution is again characterized by a non-oscillatory behavior with the magnetic energy constraint to two small section close to the endcaps (see Figure 7b), indicating that we are in the low efficiency branch.
III.4 Wall layer plus outer layer
@@
The high-efficient branch can also be achieved by assuming a configuration, where the container wall is emulated by a thin co-rotating layer with , while the exterior domain is modeled by a thick resting outer layer with . Here both additional layers have the same diffusivity, i.e. (actually, since the thick outer layer emulates the exterior this should be an insulator with ). In this case, we again obtain an oscillating solution that occupies the entire cylinder volume (see Figure 8) and has a similar structure as in the case of the thin layer or pseudo-vacuum boundary conditions without an outer layer shown in Figures 4 and 5. However, the critical magnetic Reynolds number of is now about two and a half times as large as in the previous cases and the amplitude of the oscillating part remains rather small.
III.5 Impact of the structure of the velocity field
In the following we focus on the setup with wall layer and outer layer for which we applied various time averaged velocity fields obtained within the interval . It is known from the hydrodynamic investigations that the structure of the velocity field changes only slightly within the transition between subcritical and supercritical state. Nevertheless, these gradual changes are sufficient to cause a variation in the critical magnetic Reynolds number, especially at the borders of the transition region. Figure 9 shows the growth rates for the model with wall and outer layers but varying the flow structure by using time-averaged flow fields obtained at different . Similar to the results with wall layer only (Figure 6), we see a behavior for some velocity fields that can be described by an inverse parabola, at least around the range with positive growth rates. This does not apply to velocity fields with or with . For these parameters, the parabolic behavior does not occur, and positive growth rates are found - if at all - for very large .
The intersection of the growth rates with the abscissa reveals the critical magnetic Reynolds number required for the onset of a dynamo. Figure 10 shows this number as a function of for two different setups. The blue curve shows for the original setup without any outer layer. The most outstanding feature of the blue curve is the reduction of the critical magnetic Reynolds number in the range with the lowest value being , obtained for the flow field from hydrodynamic simulations at . Below no more dynamos were found, while above the critical magnetic Reynolds number is in the range of about . The region coincides exactly with the transition region from the subcritical to the supercritical state of the velocity field, and we therefore assume that the axially symmetric double roll occurring exclusively in this parameter range plays a significant role for the onset of dynamo action at comparatively small magnetic Reynolds numbers (see Giesecke et al. [34]).
This behaviour changes, when considering the case with two outer layers that is represented by the red curve in Figure 10. In these models the onset for dynamo action is shifted to larger magnetic Reynolds numbers so that the minimum is now which is found for the flow field obtained at . Furthermore, the range of flow fields with a (relatively) small critical magnetic Reynolds number narrows and only spans the interval . Interestingly, although we still use the same structure for the flow fields, we get a large even when the double roll mode is present with a significant amplitude. This indicates that the impact of the double roll mode might be less significant than expected from the cases without outer layers but that the electromagnetic coupling of the field induced essentially by shear close to the fluid boundary layers with the side walls with finite electrical conductivity has a similarly large influence.
IV Full set of magnetohydrodynamic equations
In the following we present results obtained from the simulation of the induction equation coupled with the Navier-Stokes equation in a self-consistent MHD model.
IV.1 Code scheme
@@
The full magnetohydrodynamic problem additionally requires a computation of the temporal evolution of the flow field. In this case, the Navier-Stokes equation needs to be solved as well, which for a precessing flow in the rotating frame of reference takes the form
@@
In addition to the terms commonly encountered in regular flow problems describing nonlinear interactions and dissipation, we have three additional contributions: the Coriolis force due to rotation, the Lorentz force due to the interaction with the magnetic field, and the Poincaré force accounting for acceleration due to perpetual change of the orientation of the cylinder’s axis of rotation. A detailed derivation of the Navier-Stokes equation for a precessing system is given e.g. in Tilgner [28].
@@
We conduct direct numerical simulations (DNS) of the coupled system of equations (1) and (2) using the code SpecDyn[13, 44], which is based on a pseudo-spectral Fourier approach. The drawback of this approach is the requirement of periodic boundary conditions in all three directions. Since this rules out an application of a closed cylindrical geometry, the scheme is complemented by a direct forcing immersed boundary method (IBM) which allows the consideration of nearly arbitrary shaped boundaries that prescribe the confinement for the fluid flow when the boundary conditions on these boundaries are of no-slip type. The fulfilment of the divergence-free condition for the velocity field is ensured using a Pressure-Corrector method. The same scheme is also applied to the magnetic field to ensure that the solution of the induction equation is divergence-free as required by Maxwell’s equations. The code is parallelized using a two-dimensional processor grid. To apply the Fast Fourier Transform (FFT) to the distributed data along all spatial directions, the processor grid must be transposed, which is achieved by leveraging advanced MPI functions to perform an all-to-all communication of the processor data. Tests have been conducted with various configurations that proved accurate solutions with excellent convergence properties and nearly perfect strong scaling properties. A detailed description is given in Wilbert, Giesecke, and Grauer [13] and in Wilbert [44] and a comparable approach was used for numerical models of the VKS dynamo[45, 46].
IV.2 Model set-up
@@
We consider a setup comparable to the kinematic models with a stagnant outer layer, i.e. a setup given by a cylindrical fluid region with radius and height embedded in a Cartesian domain with a edge length so that the size of the box provides sufficient space to let the magnetic field decline outside the cylindrical vessel. The cylindrical vessel is implemented by imposing internal conditions on the fluid that emulate the (virtual) container walls by enforcing no-slip conditions at the surface of the container. Unlike in the kinematic (multi-)layer models we do not model any special region representing the wall of the vessel because such a layer would not be properly resolved. Instead we use a magnetic diffusivity for the fluid region, whereas the ’non-liquid’ region outside the vessel has a magnetic diffusivity , which is supposed to emulate an electrically isolating exterior. All simulations are carried out with a Reynolds number which is the maximum possible at the resolution of while still sufficiently resolving the boundary layers at the fluid-wall interface of the vessel [44].
IV.3 Results
@@
Similar to the procedure in the kinematic case, we vary the Poincaré number , whereby different magnetic Reynolds numbers are considered, beginning with high values. If a positive mean growth rate is found for a particular set of parameters, further runs are executed consecutively lowering the value of until no more dynamo action can be observed. Similar to the kinematic models we find a limited range for the preferred occurrence of a dynamo in the interval . However, the critical magnetic Reynolds number is significantly higher with a minimum value of at . The slightly larger value for the optimum compared to the kinematic models can be explained by the (slightly) smaller (see also Giesecke et al. [14]). In the following we focus on the particular case and and compare the properties of this DNS solution with the results found in the kinematic case. The chosen paradigmatic example well characterizes the DNS solutions found very close to the threshold while allowing a faster growth with shorter time period until a statistically steady state is reached.
IV.3.1 Temporal evolution of the kinetic energy
@@
The temporal evolutions of kinetic and magnetic energy and are displayed in Fig. 11, where and are defined as
and where denotes the edge length of the cubic domain. The simulations start with the pure hydrodynamic case, which is processed until a quasi-stationary statistically steady state is reached. A detailed description of the hydrodynamic behavior is given in Wilbert, Giesecke, and Grauer [13]. Here we focus on the properties of the magnetic field in the dynamo case and the resulting feedback on the velocity field. After switching on the magnetic induction at , the magnetic energy begins to grow for about cylinder rotations and finally goes into saturation.
@@
Both, kinetic and magnetic energy are dominated by contributions of the azimuthal components of the related field. However, the dominance for the magnetic field is not particularly pronounced and the induced magnetic energy remains much smaller than . We find that, even in the saturated state, the presence of the magnetic field does not substantially alter the structure and the amplitude of the fluid flow via the Lorentz force. A striking property of the magnetic energy is that, unlike the kinetic energy, which remains relatively steady, the magnetic energy is characterized by strong occasional peaks where suddenly rises to times above its average value. There are no related substantial changes in the flow field due to the back-reaction of the magnetic field so that the bursts can be clearly assigned to corresponding peaks in the interaction parameter , which compares the amplitude of the Lorentz force to the inertial term in the Navier-Stokes equation:
For a self-excited dynamo follows (see Miralles, Plihon, and Pinton [48]), which is indeed the case as shown in Figure 12, where we see similar irregular peaks due to the bursts shown by the magnetic energy in Figure 11b. However, the impact of the magnetic field remains small even in the case of a burst.
In the following, we discuss three characteristic periods as labeled by capital letters A, B, (two paradigmatic peaks) and C (quiescent period) in Figure 12. Figure 13 shows three rows each with four snapshots of the spatial distribution of the magnetic energy taken at four arbitrary times within each related period. Animated sequences showing the time development within the marked periods can be found in the supplementary material of this study[47]. We recognize a small scale field formed by flux patches distributed randomly in space and time. These patches are short-lived despite the strong concentration of magnetic energy, and even if they appear regularly, no corresponding period can be assigned. The magnetic field patches are slightly elongated in the azimuthal direction and during a typical peak period, we see an increased excitation of medium-scale concentrations of magnetic energy in terms of elongated arc-like magnetic structures near the end caps (Figure 13a–13h). However, these mid-scale magnetic field structures remain isolated and decay within 1-2 rotation periods, whereby the resulting small scale patches have a rather isotropic character and are distributed within the entire volume of the cylinder.
@@
While the distribution of the magnetic energy in individual snapshots appears to be largely chaotic (with the exception of the bursts), this is no longer the case when we look at the time average in the turntable system. Applying a long averaging time of more than 200 rotation periods, we obtain a rather regular distribution of magnetic energy as shown in Figure 14. Similar to the low-efficient branch of the kinematic models we find elongated structures close to the end caps, which are a result of the statistical distribution of the magnetic field energy.
Furthermore, and in contrast to the kinematic models, we recognize different symmetries with respect to the mid-plane of the cylinder, which are expressed in the different signs for the time-averaged radial field. We therefore see both, a dipole-like state and a quadrupole-like state, which are alternately dominant at irregular intervals. Due to the small-scale structure, however, the characterization as a dipole or quadrupole field is not really justified. In this sense, the results are consistent with other dynamo simulations in which a mechanical forcing was utilized as driving mechanism (see e.g. Lin et al. [49], Cébron et al. [5], Reddy, Favier, and Le Bars [50]). In agreement with Landeau’s conclusions[9], we therefore confirm that it remains unclear whether mechanical forcing can generate a large-scale magnetic field, and apart from experimental investigation, only simulations with massively reduced viscosity (i.e. massively larger Reynolds number) can help to clarify this question, which is currently prevented by insufficient performance of numerical algorithms and HPC systems.
IV.3.2 Back-reaction and kinetic power spectrum
In contrast to the magnetic field, the velocity field is always dominated by large scales as shown in Figure 15 by means of a paradigmatic time-snapshot of the three-dimensional distribution of the axial velocity component . This figure shows the typical structure of the velocity, which looks very similar at any time, regardless of whether with or without a magnetic field, or whether within a peak period or within a magnetically quiet section.
Indeed, a reduction of the speed amplitude and/or of the field structure is not absolutely necessary, because here we deal with a small-scale dynamo for which theoretical considerations (in a mean-field model) have shown that the feedback via the Lorentz force can be described by an increased diffusivity or hyperdiffusivity (Subramanian [51]).
A response of the fluid flow can be seen in a subtle way when looking at the temporal power-spectra of kinetic energy. Figure 16 compares the temporal power spectrum of the kinetic energy for the hydrodynamic case (left panel) with the spectrum for the dynamo case (right panel). Here we assume that in the high frequency range the kinetic energy of the individual components and scales according where is the energy from the Fourier component with the frequency and is a spectral index that turns out to be different for the horizontal coordinates () and the vertical coordinate (). In the hydrodynamic case, the contributions to high frequencies by the axial component (spectral index ) drop significantly faster than the contributions of the horizontal components and (spectral index ). In the dynamo case, the spectral index of the horizontal components approaches which is close to the axial component, which practically does not change compared to the hydrodynamic case. In summary, in the magnetic case we find an isotropization of the spectrum and a damping of the temporal fluctuations, which is expressed in the steeper spectrum.
V Summary and Conclusions
@@
We have run numerical simulations of precession driven dynamo action in a cylindrical cavity. We examined simplified kinematic models as well as a particular setup of a similar system solved by a self-consistent three-dimensional magnetohydrodynamic approach. The kinematic models reveal the existence of two different branches for dynamo action, the realization of which depends on the configuration of the outer boundary regions. The first branch is characterized by a highly localized and stationary dynamo that only occurs at rather large magnetic Reynolds number. The second type of dynamo is of oscillatory type with the magnetic field filling a larger proportion of the fluid domain. This type can occur at much lower magnetic Reynolds number, which, however requires suitable boundary conditions in terms of either a thin outer layer with large magnetic diffusivity or a comparable emulation with pseudo-vacuum BC.
We have not investigated here at which exact parameters the transition from one branch to the other takes place, since the experimental realization must in any case start from a basically infinitely large outer volume of an insulator, namely the air in the laboratory. The kinematic model with two outer layers (for the container walls and for the lab environment), which we assume best reflects the experimental setup, has a critical magnetic Reynolds number of , which agrees surprisingly well with the value derived by Goepfert and Tilgner [30] obtained from simulations in a precessing cube. Our solutions with external outer walls and an increased diffusivity (i.e. a decreased electrical conductivity) are consistent with previous studies[52, 53], who found a monotonic reduction of the dynamo threshold for increasing conductivity of an outer wall. Furthermore, these studies also reported a reduction for oscillatory solutions with an optimum at a particular thickness of the outer wall, which was explained by the increased dissipation resulting from eddy currents induced in the outer wall.
@@
We also used direct numerical simulations in order to investigate the time evolution of the velocity field, the impact of the magnetic field through the Lorentz force and the consistency with the kinematic models. Unfortunately, it is not possible to achieve the large hydrodynamic Reynolds number, as it is relevant in the experiment, due to the limitations of the numerics and the hardware, so that we find dynamo solutions in the DNS only for relatively large magnetic Reynolds numbers. This means that we effectively deal with a magnetic Prandtl number of the order of one and that the magnetic field is of small-scale type. At this stage it cannot be concluded with certainty that this behavior will change if we move to larger (hydrodynamic) Reynolds numbers, as there are currently insufficient parameter studies available for such scaling. Nevertheless, for the parameters considered, a small-scale burst dynamo is the typical solution. These bursts are characterized by peaks in magnetic energy, which represent the sudden excitation of large-scale magnetic field structures. Since these structures are isolated and not sustained throughout the entire volume, they are rapidly destroyed by magnetic diffusion. Our results align with findings obtained for example by Etchevest, Fontana, and Dmitruk [54], Richardson, Hollerbach, and Proctor [55], Hughes, Cattaneo, and Thelen [56], which conclude that increasing the magnetic Prandtl number leads to a transition from large-scale dynamos to small-scale dynamos.
@@
The small scale characteristics of the magnetic fields from the DNS are clearly different from the properties of the field obtained in the kinematic simulations. However, if the magnetic field from the DNS is averaged over a sufficiently long period of time so that the occasional occurrence of the strongly localized patches is also averaged out, the structure is comparable to the structure of the weakly efficient branch in the kinematic models, which only appears at large magnetic Reynolds numbers. This explains, at least in part, the high magnetic Reynolds number required to obtain a dynamo in the DNS.
In order to obtain a dynamo solution that can also be realized in the planned experiment, it would be necessary to find a path that leads to the efficient dynamo branch with low . Whether this exists at all in the full MHD case, and how this can be achieved, remains unknown at present.